OpenStudy (anonymous):

Solve for y. ln(y-1)-ln2=x+ln x

myininaya (myininaya):

ln(y-1)-ln(2)-lnx=x use this: ln(a)-ln(b)-ln(c)=ln(a)-[ln(b)+ln(c)]=ln(a)-ln(bc)=ln[a/(bc)] and write x as lne^x

myininaya (myininaya):

so then you would have \[\frac{a}{bc}=e^x\]

OpenStudy (anonymous):

Uhmm.. e^(ln(y-1/2)) = e^(x+ln x) ??

myininaya (myininaya):

no i wrote everything with a ln on one side did you what i wrote above?

OpenStudy (anonymous):

\[y=1+e^{x+\ln(2x)}\]

myininaya (myininaya):

\[\frac{y-1}{2x}=e^x\]

myininaya (myininaya):

\[y-1=2xe^x\] \[y=2xe^x+1\]

OpenStudy (anonymous):

Oh okay I see now! So log and ln have the same properties?

myininaya (myininaya):

yes they are both log \[\ln(x)=\log_e(x)\]

myininaya (myininaya):

that subscript is tiny but its suppose to be an e

OpenStudy (anonymous):

Yes that is an e

myininaya (myininaya):

ok i couldn't read my own typing